OpenStudy (anonymous):

PLEASE HELPPPPPPPPPPPPPPPPPPPPP MY MIND IS ABOUT TO BLOW UP !!! A certain dealership has a total of 100 vehicles consisting of cars and trucks. 1/2 of the cares are used, and 1/3 of the trucks are used. If there are 42 used vehicles altogether, how many trucks are there?

OpenStudy (anonymous):

half of the cars*

OpenStudy (anonymous):

i know the answer i just need to know how to get the answer

OpenStudy (anonymous):

make them into a series of two equations

OpenStudy (anonymous):

100= c+t 42 = 1/2c + 1/3t

OpenStudy (anonymous):

can you solve that?

OpenStudy (anonymous):

whoa nice..... lemme write that down

OpenStudy (cj49):

let the number of cars be x therefore trucks = x-100 x/2+(100-x/3)=42 3x+200-2x/6=42 x=42*6-200 x=252-200 x=52

OpenStudy (anonymous):

hey its not 52....

OpenStudy (anonymous):

zimmah whats the next step

OpenStudy (cj49):

x is 52 trucks=100-52=48

OpenStudy (radar):

He didn't say that the number of used cars was 52. Read the solution, he just found how many cars.......All you need to do is take half of that to find the used cars.

OpenStudy (anonymous):

make the both equations so that you can subtract them from each other to remove C. So, in that case, if you multiply the equation of 1/2c+1/3t=42 by 2, than you get c+2/3=84 (which is still the same) but now you can subtract this from the other equation. Then you get an equation that only contains t and a number. So that way, you can know t.

OpenStudy (anonymous):

can you explain to me how you cam across that equation?

OpenStudy (anonymous):

c+2/3t=84 obviously, made a typo

OpenStudy (anonymous):

1/2c+1/3t=42 you mean?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

that's the amount of used vehicles, since 1/2 of the cars is used, and 1/3 of the trucks is used. And we know that together they are 42, because that was a given.

OpenStudy (anonymous):

so what comes next?

OpenStudy (anonymous):

In most of this type of question, you really have to puzzle a bit how you can best manipulate the data to get the information you need to solve it.

OpenStudy (anonymous):

if you have two equations with both the same two unknowns (like you have here) one of the easiest methods is trying to erase one of the values from the equation. There is another way (substitution), if you think that's easier. But this way is just multiplication and subtraction. We have 100 = c+t and 42 = 1/2c + 1/3t we can easily erase the c by multiplying the second equation by 2, so that that equation also has a whole c, like the first. If you then just simply subtract the whole equation from the first, you'll have an equation that has a number on one side, and a fraction of t on the other side. That way, you know the total number of trucks. (and by extension, cars).

OpenStudy (anonymous):

hmmmmm okay can you go into the steps?

OpenStudy (anonymous):

100 = c+t 84 = c + 2/3t - --------------------- 16 = 1/3t t = 3*16 = 48

OpenStudy (anonymous):

notice how the c cancels itself out

OpenStudy (anonymous):

also, since the first equation states 100 = c+t and you now know t, you also immediatly know that c must therefore be 52 (because 100-48 = 52)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how did u get that 84?

OpenStudy (anonymous):

and whered the 2/3 come from?

OpenStudy (anonymous):

well, we had two equations to start with right?

OpenStudy (anonymous):

And both of them have not 1, but two unknown values. Which on their own, makes them impossible to solve.

OpenStudy (anonymous):

but together, you can use them to erase one of the unknown values (in this case i chose c, because that was the most logical thing to do in this case, but i might as well have chosen to erase t instead).

OpenStudy (anonymous):

But since you have to either subtract or add (sometimes you have a negative value in one of the equations, then you must add instead). It will only really erase something if one of the values is the same. Therefore i had to multiply the whole equation by 2, before subtracting.

OpenStudy (anonymous):

which equation is that?

OpenStudy (anonymous):

If used this instead 100 = c+t 42 = 1/2c + 1/3t - --------------------- 58 = 1/2c+1/3t then i would not be any better off

OpenStudy (anonymous):

because then i still know nothing about c or t

OpenStudy (anonymous):

ohhhh okay

OpenStudy (anonymous):

for that reason, i multiplied 42 = 1/2c + 1/3t by 2 so 2*42 = 2*(1/2c) + 2*(1/3t)

OpenStudy (anonymous):

What i did was \[42/\frac{ 1 }{ 3 }=126\]

OpenStudy (anonymous):

then i divided 126/.5 and I got 252

OpenStudy (anonymous):

is that wrong?

OpenStudy (anonymous):

yes, that's wrong because 42 is the total number of used vehicles, not the number of used trucks.

OpenStudy (anonymous):

so you can't just multiply 42 by 3

OpenStudy (anonymous):

and especially not multiply it again by 2. Your number is much too big

OpenStudy (anonymous):

i divided

OpenStudy (anonymous):

dividing by a fraction is the same as multiplying by the opposite of that fraction. try it out, 42*3 = 126 while 42/\(\frac{1}{3}\)=126 as well

OpenStudy (anonymous):

well why did you multiply it by 2?

OpenStudy (anonymous):

how else would you get rid of the c?

OpenStudy (anonymous):

i dont understand that process.... how do u get rid of the c by multiplying it by 2?

OpenStudy (anonymous):

you have two equations 100= c+t and 42 = 1/2c + 1/3t in order to solve them, you have to get rid of either c or t.

OpenStudy (anonymous):

because c - c = 0 but c - 1/2 c = 1/2c

OpenStudy (anonymous):

since the second equation has 1/2c, we have to multiply it by 2, otherwise we can't get c-c=0

OpenStudy (anonymous):

There is also another way, maybe that one is more your style.

OpenStudy (anonymous):

yea this one is a bit confusing....

OpenStudy (anonymous):

i wish i understood where u got that 2 from

OpenStudy (anonymous):

we have 100= c+t and 42 = 1/2c + 1/3t so to get rid of c this time we will express c as a function of t. so we move the c in 100=c+t to one side, and the rest to the other so c = 100-t (it's the same as 100=c+t, just in a different order)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

knowing that, we can use that in the second equation. 1/2*c + 1/3t = 42 since the other equation tells us c = 100-t, we can switch c out for 100-t, so let's do that. don't forget to use brackets! 1/2*(100-t) + 1/3t = 42 now let's solve the brackets first 1/2*100 = 50 and 1/2*t is just 1/2t so 50-1/2t+1/3t = 42

OpenStudy (anonymous):

i think you can solve t from there

OpenStudy (anonymous):

workin on it now

OpenStudy (anonymous):

it's a bit more work than the method i used, but at least (hopefully) you can see what is happening here.

OpenStudy (anonymous):

do i divide?

OpenStudy (cj49):

lol he is doing the samethning which i did

OpenStudy (anonymous):

42 divided by .5?

OpenStudy (anonymous):

sorry im so slow at this

OpenStudy (anonymous):

how did you arrive at 42 divided by 0.5 exactly?

OpenStudy (cj49):

hey u there?

OpenStudy (anonymous):

yes i am i was lookin over the steps on top

OpenStudy (anonymous):

if you want to solve 50-1/2t+1/3t = 42 probably best to first make 1/2t+1/3t into one fraction

OpenStudy (cj49):

lets start from the beginning let the total number of cars be x

OpenStudy (anonymous):

total number of ars is 100 no?

OpenStudy (anonymous):

please don't confuse the two methods with each other, they are both correct but very different from each other. Use only the second one, because the first one is just not for you apparently. Nothing to be ashamed of, just don't use it.

OpenStudy (anonymous):

vehicles*

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

cj you take this over? I'm helping someone else right now

OpenStudy (anonymous):

THANK YOU ZIMMAH

OpenStudy (cj49):

understand the question first we have total 100 vehicles

OpenStudy (anonymous):

yea

OpenStudy (cj49):

so lets say out of 100 we have x number of cars

OpenStudy (anonymous):

yea

OpenStudy (cj49):

so trucks wud be 100-x

OpenStudy (anonymous):

yea

OpenStudy (cj49):

now the next step in the question its mentioned that 1/2 of the cars are used

OpenStudy (anonymous):

yes

OpenStudy (cj49):

so 1/2 (x)=x/2

OpenStudy (anonymous):

why is it x/2?

OpenStudy (cj49):

bc 1/2 cars are used and we termed x as the number of cars. so out of x 1/2 cars are used. now multiply 1/2 by x and ull get the ans as x/2

OpenStudy (anonymous):

oh hmmm okay

OpenStudy (cj49):

u understood this much ryt

OpenStudy (anonymous):

yes

OpenStudy (cj49):

now we have 100-x total amt of trucks and out of that 1/3 trucks are used so now wat wud be the equation foh trucks

OpenStudy (anonymous):

x/3?

OpenStudy (cj49):

no...we have 100-x no of trucks so we wud multiply it with 1/3

OpenStudy (anonymous):

x/3 + x/2

OpenStudy (cj49):

wat is x in our equation

OpenStudy (anonymous):

the number of used trucks

OpenStudy (cj49):

scroll up ull see x=total number of cars 100-x=total number of trucks

OpenStudy (anonymous):

oh we're looking for cars? okay

OpenStudy (cj49):

no we are looking for trucks

OpenStudy (cj49):

read the question.